- What is the mass of 0.25 moles of Na2O?
Answer #1

Mass of 1 mole Na2O = 2×Na + O = 2×23+16 = 46+23 = 62 g

Mass of 0.25 moles of Na2O = 0.25×62 g

= 15.5 g Na2O

Please give feedback!